OpenStudy (anonymous):

missed 9 questions please help my test is timed

OpenStudy (anonymous):

Wind farms are a source of renewable energy found around the world. The power P (in kilowatts) generated by a wind turbine varies directly as the cube of the wind speed v (in meters per second). If a turbine generates 700kW in a 12 m/s wind, how much power does it generate in a 6 m/s wind? A. 1050kW B. 175kW C. 175kW D. 88kW

OpenStudy (anonymous):

Divide -3x3 - 2x2 - x - 2 by x – 2. A. -3x2 + 4x + 15, R 32 B. -3x2 + 4x + 15 C. -3x2 - 8x - 17 D. -3x2 - 8x - 17, R –36

OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67) A. y = -x3 + 4x2 + 2x - 2 B. y = x3 + 4x2 + 2x - 2 C. y = x4 + 4x3 + 2x2 - 2x D. y = x3 - 4x2 - 2x - 2

OpenStudy (anonymous):

Use synthetic division to find P(–2) for P(x) = x4 + 9x3 - 9x + 2 . A. –2 B. 0 C. –36 D. 68

OpenStudy (anonymous):

Divide 4x3 + 2x2 + 3x + 4 by x + 4. A. 4x2 - 14x + 59 B. 4x2 + 18x - 53, R 240 C. 4x2 - 14x + 59, R -232 D. 4x2 + 18x - 53

OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x3 + x2 - 11 A. The relative maximum is at (–1.53, 8.3) and the relative minimum is at (1.2, –12.01). B. The relative maximum is at (–1.53, 12.01) and the relative minimum is at (1.2, –8.3). C. The relative maximum is at (–1.2, 8.3) and the relative minimum is at (1.53, –12.01). D. The relative maximum is at (–1.2, 12.01) and the relative minimum is at (1.53, –8.3).

OpenStudy (anonymous):

Wind farms are a source of renewable energy found around the world. The power P (in kilowatts) generated by a wind turbine varies directly as the cube of the wind speed v (in meters per second). If a turbine generates 700kW in a 12 m/s wind, how much power does it generate in a 6 m/s wind? A. 1050kW B. 175kW C. 175kW D. 88kW

OpenStudy (anonymous):

What is the fourth term of (d + 4b)3? A. -b3 B. -64b3 C. b3 D. 64b3